2. (20 points) Costa Rica and Honduras produce two goods, bananas and pineapples. The leaders of the two countries got togetha,b,c
Answer #1

Answer:

a. From the given table we know that output per worker for banana is 10 in Costa Rica and 6 in Honduras. As Costa Rica has higher output per worker, Costa Rica has the absolute advantage in the production of bananas.

Similarly, in the production of pineapples, Costa Rica has the absolute advantage as each worker in Costa Rica can produce 6 pineapples while output per worker in Honduras is 1 pineapple. So Costa Rica has absolute advantage in the production of pineapple too.

b.Production possibility frontier for Costa rica and Honduras are as follows:

180 MANA Costa Rica Pineapple   AROLO Honduras Pineapple.   

If we draft the PPF for both the countries on the same graph, it would look like this:

  Во текол TO Costa Rica Honduras Pineapple

c. Opportunity cost (OC) for Costa Rica is as follows:

OC of Banana = 5 pineapples per 10 banana = 5/ 10 = ½

OC of pineapple = 10 bananas for 5 pineapples = 10/ 5 = 2

Opportunity cost (OC) for Honduras is as follows:

OC of banana = 1 pineapple per 6 bananas = 1/6

OC of pineapple = 6 bananas for 1 pineapple = 6/1 = 6

OC in the tabular form for both the countries:   

Costa Rica Honduras
Banana 1/2 1/6
Pineapple 2 6